OpenStudy (anonymous):

im so bad at fractions :( 1/2 + 1/3 + 1/6 this always gets me mixed up

OpenStudy (eujc21):

find the common denominator

OpenStudy (anonymous):

how would I do that? :D

OpenStudy (eujc21):

Solve for \[\frac{3}{3} \frac{1}{2} + \frac{2}{3} \frac{1}{3} + \frac{1}{6}\]

OpenStudy (eujc21):

3, 2,6 all have 6 in common.

OpenStudy (anonymous):

so the bottom number would have to be six and the top three?

OpenStudy (eujc21):

well as you can see I set up 3/3 time 1/2.... when you multiply across the fraction become 3/6

OpenStudy (eujc21):

you can do this because 3/3 = 1... so it doesn't affect the outcome/

OpenStudy (anonymous):

thankyou! :D

OpenStudy (eujc21):

3/6 is not the final answer.

OpenStudy (anonymous):

ohhh ok

OpenStudy (eujc21):

do you understand what I man when I say to multiply across.

OpenStudy (anonymous):

yes so I multiply 2x3x6 ?

OpenStudy (eujc21):

oops

OpenStudy (eujc21):

(3/3 * 1/2) + (2/2 * 1/3) + 1/6

OpenStudy (eujc21):

3/6 + 2/6 +1/6 = .....

OpenStudy (eujc21):

add the numerators

OpenStudy (anonymous):

6/6?

OpenStudy (eujc21):

yeap

OpenStudy (anonymous):

thanks :D such a big help